# Few-Shot Adversarial Low-Rank Fine-Tuning of Vision-Language Models

### v2

@@ 5.3 Practical Guidance on Selecting the Hyperparameter $\tau$
Figure 4: Average performance of AdvCLIP-LoRA with ViT-B/16 backbone on various datasets, showing clean accuracy and robust accuracy for FGSM and PGD attacks with different values of $\tau$.
Figure 5: Robust accuracy of AdvCLIP-LoRA with ViT-B/16 backbone on ImageNet with different $\tau$ and $\epsilon$ values.
5.4 Comparative Analysis of AdvCLIP-LoRA and CLIP-LoRA
Table 2 presents the experimental results of CLIP-LoRA and AdvCLIP-LoRA with varying values of $\tau$, using the ViT-B/16 backbone. Our findings show that AdvCLIP-LoRA significantly enhances model robustness, increasing FGSM accuracy for a minimum of $11.04\%$ and a maximum of $42.97\%$, and PGD accuracy for a minimum of $15.67\%$ and a maximum of $62.25\%$, averaged across all datasets. Specifically, for $\tau=1$, the model demonstrates improved robustness without a significant impact on clean accuracy (the difference in clean accuracy is less than $22.58\%$ for 1 shot and less than $2.24\%$ for 16 shots, on average). As $\tau$ increases, robustness continues to improve; however, this comes at the cost of a slight decrease in clean accuracy. This effect is less prominent for larger shots. It is noteworthy that with $16$ shots, the clean accuracy decreases by an average of only $2.24\%$, while we observe a minimum improvement of $24.55\%$ in the FGSM robustness and $29.00\%$ in the PGD robustness. For clearer comparison, we visualize clean and PGD-robust accuracies for both 4-shot and 16-shot settings across ViT-B/16 and ViT-B/32 backbones in Fig. 3.
Fig. 4 shows the average performance of AdvCLIP-LoRA with ViT-B/16 across the eight datasets. The results indicate that increasing the number of shots leads to more consistent clean accuracy while yielding disproportionately larger improvements in robust accuracy. Additionally, Fig. 5 illustrates the effect of $\epsilon$ in the PGD attack. As expected, larger values of $\epsilon$ lead to a reduced robust accuracy on ImageNet. Results for more datasets are provided in Fig. 6 in Appendix A. Further experiments and analysis using the ViT-B/32 backbone are also included in Appendix A.
6 Conclusion
In this work, we presented AdvCLIP-LoRA, the first method for enhancing adversarial robustness in CLIP models fine-tuned with LoRA in few-shot settings. By formulating adversarial fine-tuning as a minimax optimization problem, we introduced a theoretically grounded algorithm that provably converges under nonconvex–strong-concavity and smoothness assumptions. Extensive empirical evaluations across eight datasets and two CLIP backbones (ViT-B/16 and ViT-B/32) demonstrate that AdvCLIP-LoRA significantly improves robustness against FGSM and PGD attacks, with minimal impact on clean accuracy. Our results highlight the practical feasibility of integrating adversarial robustness into PEFT frameworks for VLMs, opening new directions for adaptation of large pretrained models in resource-constrained and safety-critical scenarios.
Appendix A Experiment Details
Table 3: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone. Top-1 accuracy averaged over 3 random seeds is reported. Highest value is highlighted in bold.
Average ImageNet Caltech DTD Food Pets Flowers UCF SUN
1 C-AVP [37] 43.62 17.94 46.60 11.07 85.73 50.33 26.97 12.93 24.43 5.23 57.60 22.73 63.10 29.70 3.37 0.40 41.20 11.10
APT [24] 59.07 5.08 49.30 1.30 84.77 26.90 41.67 3.83 56.57 0.83 70.23 0.60 61.97 2.10 54.50 3.87 53.53 1.23
AdvPT [25] 29.96 1.44 20.17 0.43 62.97 7.60 16.73 2.60 13.27 0.00 37.93 0.13 33.97 0.43 27.03 0.00 27.57 0.37
AdvMaPLe [30] 33.52 11.38 49.27 14.60 85.53 48.37 13.63 2.93 5.27 0.30 30.67 4.97 1.40 0.10 32.70 7.07 49.70 12.67
AdvVLP [34] 32.82 11.78 49.00 15.53 85.43 48.47 15.97 4.77 1.07 0.77 29.63 3.83 19.77 6.57 11.83 1.73 49.83 12.60
FAP [34] 40.14 10.22 49.90 15.40 83.53 41.13 18.40 2.40 31.67 1.43 49.23 3.47 10.40 0.53 28.50 2.43 49.53 14.93
AdvCLIP-LoRA1 67.18 23.9 50.86 15.89 90.79 59.31 47.70 14.24 74.22 11.81 85.80 29.14 64.27 18.64 63.73 17.31 60.05 24.90
AdvCLIP-LoRA2 66.12 24.24 51.94 17.19 87.42 58.22 48.35 15.31 73.34 11.94 83.18 29.30 63.74 19.04 61.70 16.92 59.30 25.99
AdvCLIP-LoRA5 64.22 26.32 50.29 17.53 85.56 60.93 47.75 16.96 72.82 15.01 79.39 32.92 60.94 19.57 58.90 19.83 58.10 27.78
AdvCLIP-LoRA10 61.70 27.12 47.57 17.66 83.61 60.41 47.22 16.43 69.66 18.10 74.93 32.71 60.13 22.05 53.85 21.02 56.62 28.61
AdvCLIP-LoRA25 53.32 26.28 36.96 14.74 79.07 61.46 45.74 18.79 55.32 17.59 62.77 30.36 49.86 20.30 45.33 19.67 51.52 27.29
AdvCLIP-LoRA50 34.78 19.41 27.66 10.44 68.15 57.69 39.83 18.91 17.17 4.47 29.38 14.80 15.14 8.57 36.06 17.71 44.88 22.71
AdvCLIP-LoRA75 27.75 16.53 22.18 8.26 64.54 54.52 36.23 20.15 10.20 2.67 17.66 9.27 3.74 2.80 29.08 14.17 38.39 20.43
AdvCLIP-LoRA100 23.23 14.81 15.51 5.57 62.27 52.86 30.61 19.50 8.15 2.17 10.00 6.57 3.21 3.13 25.27 12.61 30.80 16.08
2 C-AVP [37] 39.24 16.23 46.23 10.90 91.25 55.23 14.27 6.93 1.05 0.10 47.13 15.10 61.47 26.93 1.73 1.07 50.77 13.57
APT [24] 63.56 5.90 48.83 1.03 89.70 31.70 45.57 4.27 60.17 0.87 72.87 1.07 67.17 3.10 65.00 3.10 59.20 2.03
AdvPT [25] 32.47 1.76 22.37 0.77 66.07 8.33 24.57 2.43 11.13 0.17 39.17 1.17 38.47 0.23 29.37 0.23 28.57 0.77
AdvMaPLe [30] 39.09 15.58 49.97 17.13 88.00 56.20 16.53 4.20 3.10 0.67 34.03 6.87 46.17 17.00 21.17 6.20 53.73 16.33
AdvVLP [34] 42.79 17.76 50.53 17.50 87.60 55.33 18.33 7.17 1.53 1.10 31.27 7.07 62.43 25.17 36.83 11.43 53.77 17.33
FAP [34] 55.18 18.14 48.53 17.83 87.73 53.90 18.40 4.33 56.90 10.53 64.23 12.67 53.10 19.57 58.50 7.03 54.07 19.30
AdvCLIP-LoRA1 72.09 28.02 59.25 20.41 91.72 64.22 53.55 19.86 76.00 15.23 84.30 29.65 78.56 25.09 67.86 21.52 65.51 28.15
AdvCLIP-LoRA2 71.84 29.85 59.27 20.88 91.32 66.17 52.78 20.57 76.27 17.05 83.57 32.95 78.20 29.56 67.62 22.05 65.73 29.58
AdvCLIP-LoRA5 71.70 33.6 58.44 21.91 91.20 68.80 53.07 22.93 75.91 21.15 84.49 39.71 77.47 35.44 67.86 26.88 65.13 31.94
AdvCLIP-LoRA10 71.15 34.8 57.44 22.92 90.99 71.16 52.96 24.29 74.93 21.99 83.48 38.62 76.74 37.72 68.57 28.05 64.07 33.68
AdvCLIP-LoRA25 68.53 36.47 52.66 21.90 90.83 73.87 52.60 28.25 72.09 22.42 83.59 43.96 71.42 40.52 64.39 27.33 60.69 33.51
AdvCLIP-LoRA35 65.92 35.82 49.24 20.08 88.60 73.79 51.54 26.83 69.83 22.63 81.82 42.85 68.01 39.99 58.71 28.07 58.63 32.31
AdvCLIP-LoRA50 61.09 33.98 42.74 16.95 86.41 73.02 49.05 26.48 57.83 18.46 80.46 40.53 62.08 38.45 53.77 26.91 56.34 31.01
AdvCLIP-LoRA75 52.39 30.60 32.63 13.46 82.68 69.82 46.99 27.72 36.20 10.63 72.20 35.62 48.36 32.97 48.61 25.17 52.45 29.44
AdvCLIP-LoRA100 37.08 24.24 16.63 7.92 80.61 67.71 45.45 29.79 12.30 3.96 38.35 23.28 19.33 14.13 42.37 22.79 41.59 24.31
4 C-AVP [37] 43.10 16.40 49.80 11.13 90.17 52.50 18.77 9.27 22.73 4.57 57.80 16.20 55.97 23.73 1.07 0.80 48.47 13.03
APT [24] 66.37 6.04 50.90 1.40 90.77 26.67 51.33 6.33 54.80 1.63 71.83 2.10 82.40 4.23 66.53 3.03 62.37 2.90
AdvPT [25] 35.32 2.07 23.40 1.33 64.97 7.30 31.70 4.37 15.23 0.37 44.13 1.73 41.97 0.63 31.17 0.47 29.97 0.40
AdvMaPLe [30] 51.01 21.61 51.27 19.00 89.53 59.40 6.43 2.40 60.00 14.83 30.70 9.03 52.20 25.37 59.73 21.30 58.23 21.53
AdvVLP [34] 55.18 23.40 51.30 19.37 89.37 59.07 22.97 10.33 41.50 11.20 67.43 18.47 51.00 25.80 59.97 21.77 57.90 21.17
FAP [34] 57.51 24.6 51.53 19.60 87.57 57.33 31.27 8.07 59.37 18.37 42.10 9.30 73.13 38.77 58.50 22.13 56.60 23.20
AdvCLIP-LoRA1 76.11 31.89 62.53 21.89 93.59 69.41 58.69 21.87 77.21 15.79 86.86 35.65 86.20 35.16 74.20 24.11 69.64 31.26
AdvCLIP-LoRA2 76.24 34.05 62.47 22.91 93.35 71.40 59.69 23.88 77.29 18.10 87.68 37.48 85.91 39.26 73.91 26.51 69.60 32.84
AdvCLIP-LoRA5 75.75 37.12 62.11 24.69 93.10 73.59 57.51 27.13 76.49 20.91 88.12 40.58 85.79 45.11 73.80 29.58 69.09 35.39
AdvCLIP-LoRA10 75.09 38.07 61.41 25.43 93.27 74.20 57.57 27.54 75.11 23.01 87.54 41.40 85.51 48.40 71.82 28.26 68.49 36.29
AdvCLIP-LoRA25 72.93 39.26 58.08 25.56 93.18 75.90 53.90 27.84 73.33 26.25 87.11 41.26 81.24 50.51 70.53 30.82 66.03 36.94
AdvCLIP-LoRA35 71.52 39.12 56.40 24.58 93.10 78.30 53.07 26.89 71.65 25.47 85.88 41.16 79.54 48.84 67.88 31.03 64.66 36.68
AdvCLIP-LoRA50 69.21 38.42 54.52 23.33 92.41 76.84 51.65 26.65 68.55 24.57 83.95 39.98 75.23 48.40 64.34 31.96 63.03 35.65
AdvCLIP-LoRA75 63.54 35.21 46.98 19.83 91.20 75.70 48.17 26.71 57.67 19.82 82.77 37.53 64.35 40.48 57.71 28.36 59.46 33.23
AdvCLIP-LoRA100 52.08 30.12 20.24 10.42 88.64 73.59 44.98 24.70 22.08 7.55 81.22 35.57 53.35 33.37 53.56 26.43 52.54 29.33
8 C-AVP [37] 43.24 17.21 46.37 11.90 91.20 50.33 23.63 11.47 1.00 0.83 57.43 17.33 55.50 23.57 18.27 4.93 52.53 17.30
APT [24] 69.76 7.56 52.03 1.80 92.37 30.83 54.43 8.70 61.57 2.33 82.87 3.10 84.00 6.00 66.53 4.30 64.30 3.40
AdvPT [25] 38.50 2.39 24.30 0.87 68.07 10.10 37.47 4.20 16.97 0.10 44.20 0.27 51.13 0.87 33.43 1.40 32.47 1.33
AdvMaPLe [30] 62.90 30.45 52.13 20.60 90.63 63.80 33.20 16.97 62.70 20.13 55.60 21.07 83.10 48.80 64.33 25.93 61.50 26.30
AdvVLP [34] 68.32 32.87 52.83 20.97 90.17 63.13 51.83 25.77 61.73 19.33 80.67 29.63 83.90 50.90 64.07 26.97 61.33 26.23
FAP [34] 67.23 34.26 52.17 21.53 89.63 62.50 52.13 25.77 61.80 23.20 79.47 34.57 81.53 52.63 60.70 26.67 60.40 27.23
AdvCLIP-LoRA1 78.82 33.28 64.97 23.16 94.04 70.06 62.06 26.42 78.07 15.99 88.91 34.37 91.35 38.29 79.33 25.75 71.83 32.18
AdvCLIP-LoRA2 78.58 35.13 64.78 24.44 94.00 72.05 61.17 22.75 78.26 19.18 88.74 36.88 90.74 44.50 79.12 26.91 71.80 34.37
AdvCLIP-LoRA5 78.01 38.13 64.53 26.49 94.16 73.91 59.75 29.49 77.61 22.26 88.83 38.27 90.50 48.92 77.24 28.60 71.44 37.11
AdvCLIP-LoRA10 77.50 39.61 63.98 27.83 93.91 74.40 59.63 31.09 76.52 24.00 88.01 39.36 90.62 52.13 76.71 30.06 70.64 38.00
AdvCLIP-LoRA25 75.18 40.24 61.98 28.66 93.55 74.81 54.31 28.61 73.77 25.47 86.45 38.48 88.71 54.53 72.98 32.01 69.65 39.36
AdvCLIP-LoRA35 73.93 39.98 60.67 27.93 93.39 75.82 54.31 28.49 71.56 25.06 86.02 39.19 86.80 52.33 70.21 32.49 68.46 38.50
AdvCLIP-LoRA50 71.75 38.55 58.75 26.69 92.86 75.29 50.89 27.30 69.10 24.62 84.46 36.79 84.08 49.05 66.88 31.22 67.00 37.43
AdvCLIP-LoRA75 69.16 35.98 56.12 24.84 91.68 73.79 49.35 25.30 65.35 21.86 82.39 33.03 81.12 46.00 63.65 28.73 63.61 34.28
AdvCLIP-LoRA100 64.88 33.85 47.53 22.94 90.47 71.32 46.63 23.88 59.80 22.16 80.46 29.65 74.06 40.11 60.53 28.87 60.54 31.84
16 C-AVP [37] 41.90 17.13 46.27 12.77 90.40 52.60 29.20 13.87 1.07 0.80 56.40 16.43 56.17 22.03 0.97 0.93 54.70 17.63
APT [24] 71.05 8.35 52.63 2.07 92.93 30.23 54.93 10.47 62.50 2.63 83.70 4.40 86.63 8.97 69.40 4.40 65.67 3.67
AdvPT [25] 40.94 2.68 24.53 1.47 68.70 9.63 43.77 5.70 18.47 0.73 46.27 0.23 56.03 0.80 36.60 0.53 33.13 2.37
AdvMaPLe [30] 71.48 38.11 52.93 21.90 92.17 68.63 57.93 32.17 65.13 25.27 83.27 36.87 87.87 58.70 68.97 31.67 63.57 29.70
AdvVLP [34] 68.76 37.01 53.23 22.10 92.37 67.97 57.53 32.73 43.30 16.50 82.93 35.57 87.70 58.70 69.10 32.80 63.90 29.70
FAP [34] 69.88 39.22 52.53 22.90 91.10 67.33 55.17 31.33 64.03 26.67 81.90 41.00 86.27 61.47 65.70 32.80 62.37 30.27
AdvCLIP-LoRA1 80.91 35.09 66.83 24.88 95.29 72.66 66.55 25.30 78.68 17.07 88.74 33.17 95.62 48.56 81.68 25.17 73.87 33.88
AdvCLIP-LoRA2 80.96 38.25 66.74 26.03 95.25 75.38 66.08 29.61 78.52 19.71 89.21 36.55 95.82 52.82 82.10 28.60 73.99 37.26
AdvCLIP-LoRA5 80.61 40.24 66.59 28.11 95.74 76.67 65.37 30.20 78.36 22.14 89.13 37.56 95.86 56.03 79.99 32.09 73.84 39.15
AdvCLIP-LoRA10 80.12 41.53 66.06 29.48 95.38 76.80 64.60 31.74 77.57 24.21 88.66 37.53 95.53 57.53 79.70 34.58 73.49 40.35
AdvCLIP-LoRA25 78.73 42.40 64.52 30.50 95.42 78.26 63.00 31.86 75.34 24.50 86.86 38.38 94.32 58.95 77.82 34.63 72.55 42.14
AdvCLIP-LoRA35 77.49 41.97 63.06 29.81 95.21 78.78 61.88 30.97 73.74 23.61 86.07 37.88 93.10 59.11 75.15 33.70 71.70 41.89
AdvCLIP-LoRA50 75.74 39.24 61.45 28.74 94.81 78.95 58.87 29.91 70.75 20.81 85.01 35.60 92.12 50.99 72.61 27.73 70.27 40.17
AdvCLIP-LoRA75 72.41 35.04 58.38 25.86 93.55 74.73 55.08 26.77 66.56 18.21 82.15 29.93 89.69 47.06 69.36 23.10 64.54 34.63
AdvCLIP-LoRA100 68.39 31.50 46.58 21.85 91.72 71.60 53.43 25.06 61.82 13.60 80.19 25.29 86.85 40.32 65.13 22.10 61.40 31.17
Figure 7: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone.
@@ Appendix A Experiment Details
Appendix B Convergence Analysis
Before presenting the main theorem, we state several key intermediate lemmas used in the proof. For notational convenience, we denote $\Phi(W:=W_{0}+BA)$ as $\Phi(BA)$, and use $\Phi(W)$ and $\Phi(BA)$ interchangeably throughout the analysis.
Lemma B.1
@@ Appendix B Convergence Analysis
Lemma B.2
Under Assumptions 4.5 and 4.6, the function $\Phi$ is $2\kappa\ell c_{B}^{2}$-smooth with respect to $A$ when $B$ is fixed, and $2\kappa\ell c_{A}^{2}$-smooth with respect to $B$ when $A$ is fixed.
@@
$\nabla_{A}\Phi(W)=\nabla_{A}f(W,\delta^{*}(W))=B^{T}\nabla_{W}f(W,\delta^{*}(W))$ $+\left(\tfrac{\mathrm{d}\delta^{*}(W)}{\mathrm{d}W}\right)^{T}\underbrace{\nabla_{\delta}f(W,\delta^{*}(W))}_{=\boldsymbol{0}}$
@@
In $(a)$ we used Assumption 4.6 and Proposition 4.7. Similarly, we can prove that $\Phi$ is $2\kappa\ell c_{A}^{2}$-smooth with respect to $B$ when $A$ is fixed. $\Box$
Lemma B.3
The iterates $\left\{A_{t},B_{t}\right\}_{t\geq 1}$ in (6) satisfy the following inequality:
@@
Proof. Using smoothness for $A$ from Lemma B.2, we can write
@@
Using smoothness for $B$ from Lemma B.2, we can write
@@
Summing (21) and (22) yields the desired inequality. $\Box$
Lemma B.4
@@
Proof. Since $f(W_{t},\cdot)$ is $\mu$-strongly concave and $\eta_{\delta}=1/\ell$, we have [36]
@@
where in $(a)$ we used (24). Since $\delta^{\star}(\cdot)$ is $\kappa$-Lipschitz, $\left\|\delta^{\star}\left(W_{t}\right)-\delta^{\star}\left(W_{t-1}\right)\right\|\leq\kappa\left\|W_{t}-W_{t-1}\right\|$. Furthermore, we have
@@
Using (26) into (25) yields the desired inequality $\Box$
Lemma B.5
@@
Combining (28) and (29) with (18) yields the desired inequality. $\Box$
Theorem B.6
Under Assumptions 4.4, 4.5, and 4.6, and letting the stepsizes be chosen as
$\eta_{w}=\Theta\left(\mathrm{min}\left\{\tfrac{1}{\kappa\ell(c_{A}^{4}+c_{B}^{4})},\tfrac{1}{\kappa^{2}\ell(c_{A}^{2}+c_{B}^{2})},\tfrac{1}{(G^{2}+\kappa\ell c_{A}^{4}c_{B}^{2})^{1/2}}\right\}\right),$ (30)
and $\eta_{\delta}=\Theta(1/\ell)$, the number of iterations required by AdvCLIP-LoRA to return an $\epsilon$-stationary point is bounded by
where $\Delta_{\Phi}=\mathbb{E}\Phi(W_{0})-\mathbb{E}\Phi(W_{T+1})$. Moreover, the gradient complexity $M$ is bounded by
@@
Proof. Performing the inequality in Lemma B.4 recursively and using $\gamma_{0}\leq D^{2}$ from Assumption 4.5 results in
@@
Combining (33) with (27), we have
@@
Summing up (34) over $t=1,2,\cdots,T+1$ and rearranging, we can write

### v3

@@ B.2 Comparative Analysis of AdvCLIP-LoRA and CLIP-LoRA
Setup. For adversarial training, we define the projection set for updating $\delta$ as an $\ell_{\infty}$-ball with a radius of $\epsilon=10/255$ across all datasets. To evaluate adversarial robustness, we implement two standard attack methods: FGSM Szegedy et al. (2013) and PGD Madry et al. (2018). For FGSM, we set $\epsilon=10/255$, while for PGD, we use $\epsilon=2/255$ with a total of 20 attack iterations.
Table 6 presents the experimental results of CLIP-LoRA and AdvCLIP-LoRA with varying values of $\tau$, using the ViT-B/16 backbone. Our findings show that AdvCLIP-LoRA significantly enhances model robustness, increasing FGSM accuracy for a minimum of $11.04\%$ and a maximum of $42.97\%$, and PGD accuracy for a minimum of $15.67\%$ and a maximum of $62.25\%$, averaged across all datasets. Specifically, for $\tau=1$, the model demonstrates improved robustness without a significant impact on clean accuracy (the difference in clean accuracy is less than $22.58\%$ for 1 shot and less than $2.24\%$ for 16 shots, on average). As $\tau$ increases, robustness continues to improve; however, this comes at the cost of a slight decrease in clean accuracy. This effect is less prominent for larger shots. It is noteworthy that with 16 shots, the clean accuracy decreases by an average of only $2.24\%$, while we observe a minimum improvement of $24.55\%$ in the FGSM robustness and $29.00\%$ in the PGD robustness. For clearer comparison, we visualize clean and PGD-robust accuracies for both 4-shot and 16-shot settings across ViT-B/16 and ViT-B/32 backbones in Fig. 5. Further results using the ViT-B/32 model can be found in Table 7.
@@
B.3 Ablation on Attack Budget $\epsilon$
B.4 Ablation on LoRA Design Choices
Table 8: Average Clean, PGD-100, and harmonic mean (HM) for LoRA variants.
Overall Average Flowers Pets SUN UCF
Method Clean PGD HM Clean PGD HM Clean PGD HM Clean PGD HM Clean PGD HM
AdvCLIP-LoRA 81.25 34.76 48.69 90.70 48.72 63.39 87.03 32.98 47.83 71.09 31.11 43.28 76.18 26.22 39.01
Vision 78.71 30.74 44.21 86.07 37.72 52.45 88.74 35.08 50.28 67.40 26.02 37.55 72.64 24.13 36.23
$W_{q}$ 80.65 30.62 44.39 87.86 37.07 52.14 88.09 33.69 48.74 70.88 28.34 40.49 75.76 23.39 35.74
$W_{v}$ 80.95 34.73 48.61 89.05 45.43 60.17 87.14 35.62 50.57 70.53 31.18 43.24 77.09 26.40 39.33
$W_{q}W_{v}$ 80.95 34.65 48.53 89.85 48.96 63.38 86.86 33.69 48.55 71.22 30.44 42.65 75.87 25.51 38.18
up 81.21 29.32 43.08 90.17 38.81 54.26 88.42 30.25 45.08 70.35 25.95 37.91 75.89 22.28 34.45
bottom 80.09 33.02 46.76 88.10 41.18 56.13 87.03 36.77 51.70 70.30 31.11 43.13 74.91 23.00 35.19
half-up 81.37 30.72 44.60 90.05 41.53 56.84 88.14 29.82 44.56 70.40 27.14 39.18 76.90 24.40 37.05
half-bottom 79.80 32.70 46.39 88.79 42.43 57.42 85.55 33.52 48.17 70.33 31.26 43.28 74.52 23.61 35.86
mid 80.45 30.98 44.73 87.82 39.95 54.92 88.31 32.38 47.39 69.92 28.40 40.39 75.73 23.18 35.50
Appendix C Convergence Analysis
Before presenting the main theorem, we state several key intermediate lemmas used in the proof. For notational convenience, we denote $\Phi(W:=W_{0}+BA)$ as $\Phi(BA)$, and use $\Phi(W)$ and $\Phi(BA)$ interchangeably throughout the analysis. Let us begin with a few definitions.
Definition C.1
A function $f$ is $L$-Lipschitz if for all $W,W^{\prime}$, we have
Definition C.2
A function $f$ is $\ell$-smooth if for all $W,W^{\prime}$, we have
Proposition C.1
Lin et al. (2020) Under Assumption 4.2, $\Phi(\cdot)$ is $2\kappa\ell$-smooth with $\nabla\Phi(\cdot)=\nabla_{W}f\left(\cdot,\delta^{\star}(\cdot)\right)$. Also, $\delta^{\star}(\cdot)$ is $\kappa$-Lipschitz.
Lemma C.1
@@ Appendix C Convergence Analysis
Using Proposition C.1 and $\|A\|_{F}\leq c_{A}$, $\|B\|_{F}\leq c_{B}$, we can prove the smoothness of $\Phi(\cdot)$ with respect to $A$ and $B$ when the other is held fixed. Formally, we state the following lemma:
Lemma C.2
Under Assumption 4.2 and boundedness of low-rank matrices, the function $\Phi$ is $2\kappa\ell c_{B}^{2}$-smooth with respect to $A$ when $B$ is fixed, and $2\kappa\ell c_{A}^{2}$-smooth with respect to $B$ when $A$ is fixed.
@@
$\nabla_{A}\Phi(W)=\nabla_{A}f(W,\delta^{*}(W))=B^{T}\nabla_{W}f(W,\delta^{*}(W))$ $+\left(\tfrac{\mathrm{d}\delta^{*}(W)}{\mathrm{d}W}\right)^{T}\underbrace{\nabla_{\delta}f(W,\delta^{*}(W))}_{=\bm{0}}$
@@
In $(a)$, we used the boundedness of the low-rank matrices and Proposition C.1. Similarly, we can prove that $\Phi$ is $2\kappa\ell c_{A}^{2}$-smooth with respect to $B$ when $A$ is fixed. $\Box$
Lemma C.3
The iterates $\left\{A_{t},B_{t}\right\}_{t\geq 1}$ in Alg. 1 (lines 8-9) satisfy the following inequality:
@@
Proof. Using smoothness for $A$ from Lemma C.2, we can write
@@
Using smoothness for $B$ from Lemma C.2, we can write
@@
Summing equation 21 and equation 22 yields the desired inequality. $\Box$
Lemma C.4
@@
Proof. Since $f(W_{t},\cdot)$ is $\mu$-strongly concave and $\eta_{\delta}=1/\ell$, we have Lin et al. (2020)
@@
where in $(a)$ we used equation 24. Since $\delta^{\star}(\cdot)$ is $\kappa$-Lipschitz, $\left\|\delta^{\star}\left(W_{t}\right)-\delta^{\star}\left(W_{t-1}\right)\right\|\leq\kappa\left\|W_{t}-W_{t-1}\right\|$. Furthermore, we have
@@
Using equation 26 into equation 25 yields the desired inequality $\Box$
Lemma C.5
@@
Combining equation 28 and equation 29 with equation 18 yields the desired inequality. $\Box$
Theorem C.1
Let Assumptions 4.1 and 4.2 hold. Moreover, assume that the low-rank matrices remain bounded by constants $c_{A}$ and $c_{B}$ in each iteration, i.e., $\|A_{t}\|_{F}\leq c_{A}$ and $\|B_{t}\|_{F}\leq c_{B}$. Then, there exists iteration $t\in\{0,\cdots,T-1\}$ for which
$\mathbb{E}\left\|\nabla\Phi\left(W_{t}\right)\right\|^{2}\leq\mathcal{O}\left(\frac{4\Delta_{\Phi}(1/\eta_{w})+\kappa\ell^{2}(c_{A}^{2}+c_{B}^{2})D^{2}}{\epsilon^{2}}\right),$ (30)
where $\eta_{w}=\Theta(\mathrm{min}\{1/\kappa\ell(c_{A}^{4}+c_{B}^{4}),1/\kappa^{2}\ell(c_{A}^{2}+c_{B}^{2}),1/(G^{2}+\kappa\ell c_{A}^{4}c_{B}^{2})^{1/2}\})$, $\eta_{\delta}=\Theta(1/\ell)$, and $\Delta_{\Phi}=\mathbb{E}\Phi(W_{0})-\mathbb{E}\Phi(W_{T+1})$. Moreover, the mini-batch size $M$ is bounded by
@@
Proof. Performing the inequality in Lemma C.4 recursively and using $\gamma_{0}\leq D^{2}$ from Assumption 4.2 results in
@@
Combining equation 32 with equation 27, we have
@@
Summing up equation 33 over $t=1,2,\cdots,T+1$ and rearranging, we can write